OpenStudy (anonymous):

Please help me with this question MEDAL AND FAN please do not answer if you don't know

OpenStudy (anonymous):

you dont even have a question !!!!

OpenStudy (anonymous):

@BioEpic @mathmath333 @AnswerMyQuestions

OpenStudy (anonymous):

k12?

OpenStudy (anonymous):

b

OpenStudy (anonymous):

No why

OpenStudy (anonymous):

B THX

OpenStudy (anonymous):

yas

OpenStudy (anonymous):

because if it was it would be a test an that's cheating

OpenStudy (anonymous):

It isn't

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{3\cfrac{3}{5}+y+5\cfrac{2}{3}=15\\~\\ \dfrac{3\times 5+3}{5}+y+\dfrac{5 \times 3+2}{3}=15\\~\\ \dfrac{18}{5}+y+\dfrac{17}{3}=15\\~\\ y=15-\dfrac{18}{5}-\dfrac{17}{3}\\~\\ =\dfrac{86}{15}\\~\\ =\dfrac{15\times 5+11}{15}}\end{align}\)

OpenStudy (mathmath333):

when 86 is divided by 15 the remainder is 11 ans the the quotient is 5

OpenStudy (mathmath333):

\(5\cfrac{11}{15}\)